26)One card is drawn from a well shuffled deck of 52 cards. is equally likely, calculate the probability that the card will be a diamond

One card is drawn from a well shuffled deck of 52 cards.

is equally likely, calculate the probability that the card will be

a diamond
Answer #1


When a card is drawn from a well shuffled deck of 52 cards, the numher of

possible outcomes is 52.

() Let A be the event 'the card drawn is a diamond'

Clearly the number of elements in set A is 13.

13 1

Therefore, P(A) =13/52=1/4

ie probability of a diamond card=1/4

answered by: Vijayan